OpenStudy (anonymous):

How many sides does an equiangular polygon have if each interior angle measures 156?

OpenStudy (anonymous):

all i have to do is find out what the angles have to add up to right?

OpenStudy (anonymous):

for an n-sided polygon, the sum of the interior angles = 180(n-2) For a regular n-sided polygon, it follows that each angle = 180(n-2)/n = 180-360/n Since you're given 156; 156=180-360/n -> -24=-360/n -> n=360/24=15

OpenStudy (anonymous):

so there would be 15 sides then.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok. thank you